# GS-ROR: 3D Gaussian Splatting for Reflective Object Relighting via SDF Priors

### v2

@@ Abstract.
3D Gaussian Splatting (3DGS) has shown a powerful capability for novel view synthesis due to its detailed expressive ability and highly efficient rendering speed. Unfortunately, creating relightable 3D assets with 3DGS is still problematic, particularly for reflective objects, as its discontinuous representation raises difficulties in constraining geometries. Inspired by previous works, the signed distance field (SDF) can serve as an effective way for geometry regularization. However, a direct incorporation between Gaussians and SDF significantly slows training. To this end, we propose GS-ROR for reflective objects relighting with 3DGS aided by SDF priors. At the core of our method is the mutual supervision of the depth and normal between deferred Gaussians and SDF, which avoids the expensive volume rendering of SDF. Thanks to this mutual supervision, the learned deferred Gaussians are well-constrained with a minimal time cost. As the Gaussians are rendered in a deferred shading mode, while the alpha-blended Gaussians are smooth, individual Gaussians may still be outliers, yielding floater artifacts. Therefore, we further introduce an SDF-aware pruning strategy to remove Gaussian outliers, which are located distant from the surface defined by SDF, avoiding the floater issue. Consequently, our method outperforms the existing Gaussian-based inverse rendering methods in terms of relighting quality. Our method also exhibits competitive relighting quality compared to NeRF-based methods with at most 25% of training time and allows rendering at 200+ frames per second on an RTX4090.
@@
submissionid: 360journal: TOGjournal: TOGccs: Computing methodologies Rendering
Figure 1. We present an SDF-aided Gaussian Splatting framework for Reflective Object Relighting (GS-ROR) from multi-view images. We show five relighting results with reflective highlights (left) and their normal estimation (right), including Potion and Tbell from NeRO (Liu et al., 2023), Helmet and Toaster from Ref-NeRF (Verbin et al., 2022), and Qilin from NeILF++ (Zhang et al., 2023), where Qilin is from the real scene. Our method demonstrates a robust geometry reconstruction for reflective surfaces and faithful material decomposition, leading to photo-realistic and real-time reflective object relighting.
@@ 1. Introduction
Creating relightable 3D assets from multi-view images has been a long-standing and challenging task in computer graphics and vision, as the decomposition of lighting, materials, and geometries is highly ill-posed. Particularly, the decomposition becomes more difficult for reflective objects, as their appearances are highly view-dependent, and a minor error on the surface leads to a significant difference. Existing approaches (Liu et al., 2023; Li et al., 2024) have shown impressive relighting quality for reflective objects by leveraging the neural radiance field (NeRF) and the signed distance field (SDF). Unfortunately, these methods require a long training and rendering time. In this paper, we aim at reflective objects relighting given multi-view images, simultaneously achieving high-quality relighting and short training/rendering time costs.
@@
Most recently, Kerbl et al. (2023) proposed 3D Gaussian Splatting (3DGS), boosting the rendering speed significantly and achieving more detailed appearance modeling. A concurrent work (Ye et al., 2024) has introduced deferred Gaussian splatting for reflective objects, significantly improving the novel view synthesis (NVS) quality. Despite the impressive NVS quality, the relighting with 3DGS becomes problematic due to the Gaussians’ discontinuity and high flexibility, especially for reflective objects. Extensive efforts have been made to improve the relighting quality by regularizing the geometry in terms of the normal and depth (Gao et al., 2024; Jiang et al., 2024; Liang et al., 2024). However, they still suffer from erroneous surfaces and floater artifacts for reflective objects.
@@
In this paper, we propose a novel framework GS-ROR for reflective objects relighting by leveraging SDF priors in the deferred Gaussian splatting. A direct incorporation between Gaussians and SDF significantly slows training, due to the expensive computational cost of SDF rendering. For this, we propose to supervise the deferred Gaussians and SDF mutually in our framework, so-called mutual-supervision, which allows constraining deferred Gaussians with SDFs while avoiding SDF rendering. Specifically, the depth/normal of deferred Gaussians is supervised by SDF for regularization, while the SDF is supervised by Gaussians via the depth/normal rather than SDF rendering. This way, the learned deferred Gaussians are better constrained with a minimal computational cost. However, we still notice that Gaussians have floater artifacts occasionally, as the deferred Gaussians (with alpha blending) rather than individual Gaussians are constrained by SDF. Therefore, we introduce an SDF-aware pruning strategy to remove Gaussian outliers to avoid this floater issue. Consequently, our method outperforms the existing Gaussian-based inverse rendering methods in terms of relighting quality. Meanwhile, it exhibits competitive relighting quality compared to NeRF-based methods with at most 25% of training time. It also allows real-time rendering of reflective objects with 200+ frames per second on an RTX4090. To summarize, our main contributions include
@@
  • a GS-ROR framework for reflective object relighting, achieving high-quality relighting while maintaining short training time and allowing real-time rendering,
@@
  • a mutual supervision approach between deferred Gaussians and the SDF, producing robust geometry and avoiding local minima with minimal computational increase, and
  • an SDF-aware Gaussian pruning strategy, which prunes outlier Gaussian, avoiding the float artifacts.
@@ 2.1. Neural representations for multi-view stereo
3DGS (2023) sheds light on the multi-view stereo in the rasterization framework, providing impressive real-time NVS results. 3DGS utilizes the discrete explicit 3D Gaussian primitive and projects these primitives onto the image plane to obtain the pixel color via alpha blending, which avoids the consuming sampling operation. However, 3DGS imposes no geometry constraint while optimizing, leading to noisy depth and normal, which hinders the usage of 3DGS in many downstream tasks. To resolve this issue, existing methods construct different constraints on Gaussians by introducing 2D Gaussians (Huang et al., 2024) or Gaussian Surfel (Dai et al., 2024). SuGaR (Guédon and Lepetit, 2024) extracts a mesh from the Gaussians and binds the flattened Gaussians to the surface to optimize further. These methods improve the geometry quality significantly, which is orthogonal to our relighting framework and can further improve our relighting quality. We leave it for future work.
Several previous works introduce SDF into 3DGS. NeuSG (Chen et al., 2023) introduces SDF into 3DGS to improve the details of surface reconstruction by optimizing NeuS and 3DGS jointly. Despite the high quality of the reconstructed surface, the training time is tens of times longer than 3DGS. GSDF (Yu et al., 2024) utilizes the depth from Gaussian to guide sampling for SDF, but they still need to render color in the SDF branch, leading to higher efficiency. While our method also incorporates the SDF and Gaussians, the key difference is that it drops SDF rendering, improving the geometry quality with a low time cost.
@@ 2.2. Inverse rendering
Inverse rendering aims to decompose geometry, material, and light from multi-view RGB images. Most existing methods adopt the geometry formulation from NeRF in either density or SDF manner, namely NeRF-based method. In terms of material modeling, most methods utilize the Disney Principled BRDF model (McAuley et al., 2013), while NeRFactor (Zhang et al., 2021b) pretrains a BRDF MLP with a measured BRDF dataset. Considering the light modeling, early works (Zhang et al., 2021a; Boss et al., 2021b; Bi et al., 2020; Boss et al., 2021a; Zhang et al., 2022a) consider direct light only, and more recent works (Srinivasan et al., 2021; Zhang et al., 2022b; Jin et al., 2023; Zhang et al., 2023; Yang et al., 2023) model indirect lighting and visibility, leading to a higher-quality decomposition. NeRF Emitter (Ling et al., 2024) incorporates NeRF as a non-distant environment emitter and uses importance sampling of NeRF to reduce rendering variance. However, these NeRF-based methods suffer from slow rendering speed and high memory usage in training.
More recently, Gaussian-based inverse rendering methods have begun to emerge. These methods regularize Gaussians to obtain better geometry in various ways. Gaussian Shader (Jiang et al., 2024) and GIR (Shi et al., 2023) link the shortest axis of Gaussian with the normal. The Gaussian shader, Relightable Gaussian (Gao et al., 2024), and GS-IR (Liang et al., 2024) supervise the Gaussian normal with the depth normal. To model the indirect illumination, Relightable Gaussian, GIR, and GS-IR compute the visibility via one-step ray tracing and represent the indirect illumination with SH coefficients. All these Gaussian-based relighting methods cannot relight reflective objects, while our method enables such scenarios.
@@ 2.3. Reflective objects NVS and relighting
Another group of methods enables relighting for reflective materials. NeRO (Liu et al., 2023) proposes a novel light representation consisting of two MLPs for direct and indirect light, respectively. With explicit incorporation of the rendering equation, NeRO reconstructs reflective objects of high quality. Wang et al. (2024) represent the indirect light with a 5-dimensional function, so-called the 5D neural plenoptic function. Thanks to material-aware integral positional encoding, their method reduces the light sample per pixel during training and provides more accurate material decomposition. TensoSDF (Li et al., 2024) designs a roughness-aware incorporation between the radiance and reflectance field for robust geometry reconstruction and proposes a novel tensorial representation with SDF, enabling fast SDF query. Although their representation is faster than the original SDF, it still needs hours of training time and cannot achieve real-time rendering. DeferredGS (Wu et al., 2024) utilizes the idea of deferred shading to improve the quality of the reflective object relighting in the Gaussian framework. However, their method needs to distill a reasonable geometry from a pretrained SDF, taking more than 3 hours and thus losing the efficiency in training. Our method takes a further step in the reflective object relighting with 3D Gaussians and provides high-quality relighting results while keeping highly efficient training and rendering.
@@
Figure 3. Overview of our framework. The architecture of our proposed method consists of two geometry representations (i.e., Gaussian primitive and TensoSDF). In the deferred Gaussian pipeline, the shading parameters (i.e., albedo a, roughness r and metallicity m), normal and depth are projected to the image plane and alpha blended. The pixel color $C_{\rm gs}$ is calculated using the split-sum approximation and supervised by ground truth color $C_{\rm gt}$. In the TensoSDF, we sample rays originated from camera center o and view direction v and query the SDF value and gradient for each point $p$ along the ray $\textbf{o}+t\textbf{v}$. The normal $\textbf{n}_{\rm sdf}$ and depth $D_{\rm sdf}$ are obtained via volume rendering, which is supervised mutually with the normal $\textbf{n}_{\rm gs}$ the depth $D_{\rm gs}$ from Gaussians. Note no color network is used in the SDF part, and only the geometry attributes are volume rendered.
@@ 3. Preliminaries and motivation
where $v^{m}_{k}$ and $M^{\widetilde{m}}_{k}$ represent the $k$-th vector and matrix factors of their corresponding spatial axes $m$, and $\widetilde{m}$ denotes the two axes orthogonal to $m$ e.g., $\widetilde{X}=YZ$. $\circ$ and $\oplus$ represent the element-wise multiplication and concatenation operations. $V_{p}$ is the latent vector from the tensorial encoder and then is decoded with the position $p$ by a tiny MLP $\Theta$ to get the SDF value $s$. TensoSDF backbone enables a faster convergence speed and low-cost SDF query.
@@ 4. Method
In this section, we present our framework GS-ROR for reflective object relighting. First, we present a deferred Gaussian splatting for relighting (Sec. 4.1). Then, we introduce the SDF priors into the deferred Gaussian splatting (Sec. 4.2) for geometry regularization. Finally, we propose an SDF-aware pruning strategy to refine the Gaussian distributions (Sec. 4.3).

### v3

# GS-ROR2: Bidirectional-guided 3DGS and SDF for Reflective Object Relighting and Reconstruction

@@ Abstract.
3D Gaussian Splatting (3DGS) has shown a powerful capability for novel view synthesis due to its detailed expressive ability and highly efficient rendering speed. Unfortunately, creating relightable 3D assets and reconstructing faithful geometry with 3DGS is still problematic, particularly for reflective objects, as its discontinuous representation raises difficulties in constraining geometries. In contrary, volumetric signed distance field (SDF) methods provide robust geometry reconstruction, while the expensive ray marching hinders its real-time application and slows the training. Besides, these methods struggle to capture sharp geometric details. To this end, we propose to guide 3DGS and SDF bidirectionally in a complementary manner, including an SDF-aided Gaussian splatting for efficient optimization of the relighting model and a GS-guided SDF enhancement for high-quality geometry reconstruction. At the core of our SDF-aided Gaussian splatting is the mutual supervision of the depth and normal between blended Gaussians and SDF, which avoids the expensive volume rendering of SDF. Thanks to this mutual supervision, the learned blended Gaussians are well-constrained with a minimal time cost. As the Gaussians are rendered in a deferred shading mode, the alpha-blended Gaussians are smooth, while individual Gaussians may still be outliers, yielding floater artifacts. Therefore, we introduce an SDF-aware pruning strategy to remove Gaussian outliers located distant from the surface defined by SDF, avoiding the floater issue. This way, our GS framework provides reasonable normal and achieves realistic relighting, while the mesh of truncated SDF (TSDF) fusion from depth is still problematic. Therefore, we design a GS-guided SDF refinement, which utilizes the blended normal from Gaussians to finetune SDF. Equipped with the efficient enhancement, our method can further provide high-quality meshes for reflective objects at the cost of 17% extra training time. Consequently, our method outperforms the existing Gaussian-based inverse rendering methods in terms of relighting and mesh quality. Our method also exhibits competitive relighting/mesh quality compared to NeRF-based methods with at most 25%/33% of training time and allows rendering at 200+ frames per second on an RTX4090. We will release the code.
@@
journal: TOGjournal: TOGccs: Computing methodologies Rendering
Figure 1. We present bidirectional-guided 3DGS and SDF for Reflective Object Relighting and Reconstruction (GS-ROR2) from multi-view images. We show relighting results with reflective highlights (left) and their normal (right) in the $1^{st}$ row, including Potion, Tbell from NeRO (Liu et al., 2023), Helmet, Toaster from Ref-NeRF (Verbin et al., 2022), and Qilin from NeILF++ (Zhang et al., 2023), where Qilin is a real scene. Besides, we show extracted meshes in the $2^{nd}$ row, including Bell, Horse, Teapot from NeRO and Car, Toaster from Ref-NeRF. Our method demonstrates a robust geometry reconstruction for reflective surfaces and faithful material decomposition, leading to photo-realistic reflective object relighting and high-fidelity geometry reconstruction.
@@ 1. Introduction
Creating relightable 3D assets from multi-view images has been a long-standing and challenging task in computer graphics and vision, as the decomposition of lighting, materials, and geometries is highly ill-posed. Particularly, the decomposition becomes more difficult for reflective objects, as their appearances are highly view-dependent, and a minor error on the surface leads to a significant difference. Existing approaches (Liu et al., 2023; Li et al., 2024) have shown impressive relighting quality for reflective objects by leveraging the neural radiance field (NeRF) and the signed distance field (SDF). Unfortunately, these methods require a long training and rendering time. They also struggle to capture sharp geometry details. In this paper, we aim at reflective object relighting and reconstruction given multi-view images, achieving high-quality relighting and detailed reconstruction with short training/rendering time costs.
@@
Most recently, Kerbl et al. (2023) proposed 3D Gaussian Splatting (3DGS), boosting the rendering speed significantly and achieving more detailed appearance modeling. A concurrent work (Ye et al., 2024) has introduced deferred Gaussian splatting for reflective objects, improving the novel view synthesis (NVS) quality. Despite the impressive NVS quality, relighting and geometry reconstruction with 3DGS become problematic due to the Gaussians’ discontinuity and high flexibility, especially for reflective objects. Extensive efforts have been made to improve the relighting quality by regularizing the geometry in terms of the normal and depth (Gao et al., 2024; Jiang et al., 2024; Liang et al., 2024). However, they still suffer from erroneous surfaces and floater artifacts for reflective objects.
Some prior works (e.g., NeuSG (Chen et al., 2023), GSDF (Yu et al., 2024a)) couple SDF with GS for surface reconstruction. While their work has shown impressive reconstructed quality for diffuse objects, they require extensive training time and memory cost, due to the rendering of SDF and the existence of two representations. Our method focuses on both relighting and reconstruction, thus raising more difficulties. Specifically, evaluating the rendering equation increases the training time, and additional material parameters make the memory cost issue more obvious. These issues become more severe when modeling reflective objects, as the specular highlight is sensitive to the reflective direction, which needs more Gaussians to provide detailed normal. Therefore, introducing SDF into GS for our task is not straightforward, which raises the demand for a more efficient framework.
@@
In this paper, we propose a bidirectional guidance of GS and SDF in a complementary manner, including an SDF-aided Gaussian splatting for efficient optimization of the relighting model and a GS-guided SDF enhancement for high-quality geometry reconstruction. Specifically, the SDF-aided Gaussian splatting utilizes the SDF to constrain the Gaussians with minimal computational increase. For this, we propose to supervise the deferred Gaussians and a low-resolution SDF mutually in our framework, so-called mutual supervision, which allows constraining deferred Gaussians with SDFs while avoiding SDF rendering. This way, the learned deferred Gaussians are better constrained. However, we still notice that Gaussians have floater artifacts occasionally, as the deferred Gaussians (with alpha blending) rather than individual Gaussians are constrained by SDF. Therefore, we introduce an SDF-aware pruning strategy to remove Gaussian outliers to avoid this floater issue. Therefore, our method provides reasonable shading normal and enables realistic relighting for reflective objects in the GS framework. Due to the inconsistency between the normal and depth from blended Gaussians, the meshes of TSDF fusion from depth are still problematic. Meanwhile, SDF is underfitted and only learns the coarse structure of geometry. To solve this, we further design a GS-guided SDF enhancement, which utilizes the learned normal from fixed Gaussians to finetune the upsampled SDF. This way, we avoid optimizing a full-resolution SDF and GS jointly, reducing 40% of the memory requirement. SDF benefits from the detailed expressive ability of GS and learns more fine-grained geometry details. Equipped with this design, our method achieves high-quality geometry reconstruction for reflective objects from SDF with only 17% extra overhead of training time. Consequently, our method outperforms the existing Gaussian-based inverse rendering methods in terms of relighting and mesh quality. Meanwhile, it exhibits competitive relighting/mesh quality compared to NeRF-based methods with at most 25%/33% of training time. It also allows real-time rendering with 200+ frames per second on an RTX4090. To summarize, our main contributions include
@@
  • a GS-ROR2 framework for reflective objects, achieving high-quality relighting and geometry reconstruction while maintaining short training time and allowing real-time rendering,
@@
  • an SDF-aided GS, including a mutual supervision between deferred Gaussians and the SDF with an SDF-aware Gaussian pruning strategy, producing robust geometry and avoiding local minima with minimal computational increase, and
  • a GS-guided SDF enhancement via blended normal from Gaussians, enabling high-quality geometry reconstruction.
@@ 2.1. Neural representations for multi-view stereo
3DGS (2023) sheds light on the multi-view stereo in the rasterization framework, providing impressive real-time NVS results. 3DGS utilizes the discrete explicit 3D Gaussian primitive and projects these primitives onto the image plane to obtain the pixel color via alpha blending, which avoids the consuming sampling operation. However, 3DGS imposes no geometry constraint while optimizing, leading to noisy depth and normal, which hinders the usage of 3DGS in many downstream tasks. To resolve this issue, existing methods construct different constraints on Gaussians by introducing 2D Gaussian (Huang et al., 2024) or Gaussian Surfel (Dai et al., 2024). SuGaR (Guédon and Lepetit, 2024) extracts meshes from the Gaussians and binds the flattened Gaussians to the surface to optimize further. These new primitives and regularization could improve the geometry quality, potentially benefiting the relighting task.
Several previous works introduce SDF into 3DGS. NeuSG introduces SDF into 3DGS to improve the details of surface reconstruction by optimizing NeuS and 3DGS jointly. Despite the high quality of the reconstructed surface, the training time is tens of times longer than 3DGS. GSDF utilizes the depth from Gaussian to guide sampling for SDF, but they still need to render color in the SDF branch, leading to higher efficiency. While our method also incorporates the SDF and Gaussians, the key difference is that it drops SDF rendering, improving the geometry quality with a low time cost.
@@ 2.2. Inverse rendering
Inverse rendering aims to decompose geometry, material, and light from multi-view RGB images. Most existing methods adopt the geometry formulation from NeRF in either density or SDF manner, namely NeRF-based method. In terms of material modeling, most methods utilize the Disney Principled BRDF model (McAuley et al., 2013), while NeRFactor (Zhang et al., 2021b) pretrains a BRDF MLP with a measured BRDF dataset. Considering the light modeling, early works (Zhang et al., 2021a; Boss et al., 2021b; Bi et al., 2020; Boss et al., 2021a; Zhang et al., 2022a) consider direct light only, and recent works (Srinivasan et al., 2021; Zhang et al., 2022b; Jin et al., 2023; Zhang et al., 2023; Yang et al., 2023; Sun et al., 2023) model indirect lighting and visibility, leading to a higher-quality decomposition. NeRF Emitter (Ling et al., 2024) incorporates NeRF as a non-distant emitter and uses importance sampling of NeRF to reduce rendering variance. Neural-PBIR (Sun et al., 2023) introduces the differentiable path tracing to model complex inter-reflection. However, these NeRF-based methods suffer from slow rendering speed and high memory usage in training.
More recently, Gaussian-based inverse rendering methods have begun to emerge. These methods regularize Gaussians to obtain better geometry in various ways. GShader (Jiang et al., 2024) and GIR (Shi et al., 2023) link the shortest axis of Gaussian with the normal. The GShader, R3DG (Gao et al., 2024), and GS-IR (Liang et al., 2024) supervise the Gaussian normal with the depth normal. To model the indirect illumination, R3DG, GIR, and GS-IR compute the visibility via one-step ray tracing and represent the indirect illumination with SH coefficients. These early Gaussian-based relighting methods cannot relight reflective objects, while our method enables such scenarios.
@@ 2.3. Reflective objects NVS and relighting
Another group of methods enables relighting for reflective materials. NeRO (Liu et al., 2023) proposes a novel light representation consisting of two MLPs for direct and indirect light, respectively. With explicit incorporation of the rendering equation, NeRO reconstructs reflective objects of high quality. Wang et al. (2024) represent the indirect light with a 5-dimensional function, so-called the 5D neural plenoptic function. Thanks to material-aware integral positional encoding, their method reduces the light sample per pixel during training and provides more accurate material decomposition. TensoSDF (Li et al., 2024) designs roughness-aware incorporation between the radiance and reflectance field for robust geometry reconstruction and proposes a novel tensorial representation with SDF, enabling fast SDF query. Although their representation is faster than the original SDF, it still needs hours of training time and cannot achieve real-time rendering. Recently some Gaussian-based methods enable reflective object relighting with the help of pretrained models. DeferredGS (Wu et al., 2024) utilizes the idea of deferred shading to improve the quality of the reflective-object relighting in the Gaussian framework. However, their method needs to distill reasonable geometry from a pretrained SDF, taking more than 3 hours on an RTX 4090 and thus losing the efficiency in training. GlossyGS (Lai et al., 2024) introduces the segmentation prior for a reasonable material estimation, while the segmentation model (i.e., DINOv2 (Oquab et al., 2024)) needs more than 3 days on 4 RTX A100 to finetune with custom data. Our method takes a further step in the reflective object relighting with 3D Gaussians and provides high-quality relighting results without any pretrained model, keeping highly efficient training and rendering. Furthermore, our method can provide high-quality mesh with efficient finetuning, while the existing Gaussian-based methods fail to provide high-quality meshes for reflective objects.
Figure 2. Overview of our SDF-aided Gaussian Splatting. The architecture of our proposed method consists of two geometry representations (i.e., Gaussian primitive and TensoSDF). In the deferred Gaussian pipeline, the shading parameters (i.e., albedo a, roughness r and metallicity m), normal and depth are projected to the image plane and alpha blended. The pixel color $C_{\rm gs}$ is calculated using the split-sum approximation and supervised by ground truth color $C_{\rm gt}$. In the TensoSDF, we sample rays originated from camera center o and view direction v and query the SDF value and gradient for each point $p$ along the ray $\textbf{o}+t\textbf{v}$. The normal $\textbf{n}_{\rm sdf}$ and depth $D_{\rm sdf}$ are obtained via volume rendering, which is supervised mutually with the normal $\textbf{n}_{\rm gs}$ the depth $D_{\rm gs}$ from Gaussians. Note no color network is used in the SDF part, and only the geometry attributes are volume rendered.
@@ 3. Preliminaries and motivation
where $v^{m}_{k}$ and $M^{\widetilde{m}}_{k}$ represent the $k$-th vector and matrix factors of their corresponding spatial axes $m$, and $\widetilde{m}$ denotes the two axes orthogonal to $m$ e.g., $\widetilde{X}=YZ$. $\circ$ and $\oplus$ represent the element-wise multiplication and concatenation operations. $V_{p}$ is the latent vector from the tensorial encoder and then is decoded with the position $p$ by a tiny MLP $\Theta$ to get the SDF value $s$. The TensoSDF backbone enables faster convergence speed and low-cost SDF queries. Besides, the capacity of TensoSDF can be modulated efficiently through the interpolation of the tensorial grid.
@@ 4. Method
This section presents our bidirectional-guided GS and SDF for reflective object relighting and reconstruction, named GS-ROR2. First, we present an SDF-aided GS for relighting, including a deferred GS pipeline (Sec. 4.1), an SDF prior of mutual supervision for geometry regularization (Sec. 4.2), and an SDF-aware pruning to refine the Gaussian distribution (Sec. 4.3). Finally, we present a GS-guided SDF enhancement for high-quality geometry reconstruction (Sec. 4.4).